# InkubaLM: A small language model for low-resource African languages

### v1

@@ 2.2 Small language models
Zhang et al. Zhang et al. (2024) introduced TinyLlama, a 1.1 billion parameter language model pre-trained on 1 trillion tokens. Despite its compact size, TinyLlama leverages techniques such as FlashAttention to deliver strong performance across various tasks, outperforming many models within its class. Building on this, the authors developed TinyLlama v1.1, which includes specialized versions of models tailored for math, code, and Chinese language tasks, showcasing improved results through a multi-stage pretraining process.
@@
In another approach to model efficiency, Song et al. Song et al. (2024) focused on achieving sparse activation in SLMs. They developed a new attribution metric to overcome the limitations of existing sparse activation techniques, successfully achieving an 80% sparsification ratio with minimal accuracy loss comparable to larger models.
@@ 2.3 LLMs for African languages
Recent research on LLMs for African languages focuses on several key aspects: the creation of linguistic resources, the development of adapted models, and the improvement of the performance of LLMs for these languages. Adelani et al. Adelani et al. (2024a) introduce IrokoBench, a benchmark dataset for 16 African languages, which uncovers significant performance gaps between high-resource and low-resource languages and underscores the importance of developing LLMs tailored to African languages, especially given the disparity between open and proprietary models. Ogueji et al. Ogueji et al. (2021) present a novel approach with AfriBERTa, a multilingual language model specifically trained on low-resource African languages, showing that “small data” approaches can outperform traditional models like mBERT and XLM-R without relying on high-resource languages. Building on the challenges of low-resource languages, Adejumobi et al. Joshua (2024) address the hallucination issues in LLMs, particularly GPT-3.5 turbo, when processing Yoruba by employing Retrieval-Augmented Generation (RAG) techniques, significantly enhancing the models’ accuracy and cultural relevance. Similarly, Lawal et al. Lawal et al. (2024) explore the performance of LLMs in educational settings, particularly in understanding and generating Yoruba primary education science content, and reveal a need for more targeted language-specific models due to the underperformance of existing LLMs in this context. Focusing on another low-resource language, Azime et al. Azime et al. (2024) enhance the LLAMA-2-Amharic model by integrating task-specific and generative datasets, demonstrating improved performance in various NLP tasks and contributing to the growing body of resources for low-resource languages.
@@ 5 InkubaLM
The environmental impact of training the InkubaLM-13B model was considered and measured using the machine learning impact calculator  [^7] https://mlco2.github.io/impact#compute. The training process on 8 A100 NVIDIA GPUs over 16 days resulted in an estimated carbon emission of 53.76 kg of CO2 equivalent. The training was performed on the Google Cloud Platform (GCP) in the Asia-Southeast1-C region, where the carbon footprint was carefully monitored to assess the environmental impact.

### v2

@@ Abstract
InkubaLM: A small language model for low-resource African languages
Atnafu Lambebo Tonja1,2,∗, Bonaventure F. P. Dossou1,3,∗, Jessica Ojo1,3, Jenalea Rajab1,5, Fadel Thior1, Eric Peter Wairagala1, Anuoluwapo Aremu1, Pelonomi Moiloa1, Jade Abbott1, Vukosi Marivate1,4, Benjamin Rosman1,5 1 Lelapa AI, 2 MBZUAI, 3 McGill University & Mila Quebec AI Institute, 4 DSFSI - University of Pretoria, 5 RAIL Lab - University of the Witwatersrand Equal Contributions
@@ 2.2 Small language models
Zhang et al. (2024) introduced TinyLlama, a 1.1 billion parameter language model pre-trained on 1 trillion tokens. Despite its compact size, TinyLlama leverages techniques such as FlashAttention to deliver strong performance across various tasks, outperforming many models within its class. Building on this, the authors developed TinyLlama v1.1, which includes specialized versions of models tailored for math, code, and Chinese language tasks, showcasing improved results through a multi-stage pretraining process.
@@
In another approach to model efficiency, Song et al. (2024) focused on achieving sparse activation in SLMs. They developed a new attribution metric to overcome the limitations of existing sparse activation techniques, successfully achieving an 80% sparsification ratio with minimal accuracy loss comparable to larger models.
@@ 2.3 LLMs for African languages
Recent research on LLMs for African languages focuses on several key aspects: the creation of linguistic resources, the development of adapted models, and the improvement of the performance of LLMs for these languages. Adelani et al. (2024a) introduce IrokoBench, a benchmark dataset for 16 African languages, which uncovers significant performance gaps between high-resource and low-resource languages and underscores the importance of developing LLMs tailored to African languages, especially given the disparity between open and proprietary models. ogueji-etal-2021-small present a novel approach with AfriBERTa, a multilingual language model specifically trained on low-resource African languages, showing that “small data” approaches can outperform traditional models like mBERT and XLM-R without relying on high-resource languages. Building on the challenges of low-resource languages, Joshua (2024) address the hallucination issues in LLMs, particularly GPT-3.5 turbo, when processing Yoruba by employing Retrieval-Augmented Generation (RAG) techniques, significantly enhancing the models’ accuracy and cultural relevance. Similarly, Lawal et al. (2024) explore the performance of LLMs in educational settings, particularly in understanding and generating Yoruba primary education science content, and reveal a need for more targeted language-specific models due to the underperformance of existing LLMs in this context. Focusing on another low-resource language, Azime et al. (2024) enhance the LLAMA-2-Amharic model by integrating task-specific and generative datasets, demonstrating improved performance in various NLP tasks and contributing to the growing body of resources for low-resource languages.
@@ 5 InkubaLM
The environmental impact of training the InkubaLM-0.4B model was considered and measured using the machine learning impact calculator  [^7] https://mlco2.github.io/impact#compute. The training process on 8 A100 NVIDIA GPUs over 16 days resulted in an estimated carbon emission of 53.76 kg of CO2 equivalent. The training was performed on the Google Cloud Platform (GCP) in the Asia-Southeast1-C region, where the carbon footprint was carefully monitored to assess the environmental impact.